OpenStudy (anonymous):

What is the value of t in the equation 3(2t + 5) = 5t + 25? 5 6 10 35

OpenStudy (anonymous):

distribute 3(2t+5) 6t+15= 5t+25 -15 -15 6t=5t+10 -5t -5t 1t=10

DivineSolar (divinesolar):

No participation..

OpenStudy (anonymous):

Thanks so much!!

OpenStudy (anonymous):

no problemo